OpenStudy (anonymous):

1/5(2x-15)+3/5x=1 1/4

OpenStudy (theloshua):

\[\left(\begin{matrix}1 \\ 5\end{matrix}\right)(2x-15) + \frac{ 3 }{ 5 }x = 1 \frac{ 1 }{ 4 }\]

OpenStudy (theloshua):

is that the equation?

OpenStudy (anonymous):

yes @Theloshua

OpenStudy (theloshua):

okay, first distribute 1/5

OpenStudy (theloshua):

can you do that?

OpenStudy (anonymous):

ok hold on

OpenStudy (theloshua):

kay :)

OpenStudy (anonymous):

i hate fractions, don't you?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

i would multiply both sides by \(5\) first

OpenStudy (anonymous):

you will get \[2x-15+3x=6\tfrac{1}{4}\] and now you have no fraction, just an annoying mixed number

OpenStudy (anonymous):

thanks and ur funny

OpenStudy (anonymous):

you good from there? combine like terms first, then add \(15\) to both sides, and finally divide both sides by \(5\)

OpenStudy (anonymous):

thank you